This is an example of the time independent Schrodinger equation with a ramp potential, V (x) = kx. Substituing this potential into the S.E. gives

\frac{d^{2}ψ(x)}{dx^{2}} = \frac{2m}{ℏ^{2}} [kx − E] ψ(x)